# Solving Robotics Tasks with Prior Demonstration via Exploration-Efficient Deep Reinforcement Learning

### v1

@@ IV Methods
To mitigate bootstrapping error in IBRL, one should consider using batches with accurate Q-value estimates to update the critic and policy networks. Prior studies have shown that the Q-value estimates of $Q_{\phi^{\prime}}\left(s_{t+1},a(s_{t+1})\right)$ are only reliable when $(s_{t+1},a(s_{t+1}))$ is from the same distributions as the dataset used to train $\hat{Q}(s_{t},{a}_{t})$ [29, 33]. In our critic networks update process, instead of selecting between $Q_{\phi^{\prime}}\left(s_{t+1},\mu_{\psi}(s_{t+1})\right)$ and $Q_{\phi^{\prime}}\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$, where both $\left(s_{t+1},\mu_{\psi}(s_{t+1})\right)$ and $\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ could be OOD state-action pairs. We propose to select between $Q_{\phi^{\prime}}\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ and $Q_{\phi^{\prime}}\left(s^{\prime}_{t+1},\mu_{\psi}(s^{\prime}_{t+1})\right)$, where $s^{\prime}_{t+1}$ are only sampled from $\mathcal{D}$. This modification ensures $(s^{\prime}_{t+1},\mu_{\psi}(s^{\prime}_{t+1}))$ is always from the same distribution as the $\mathcal{D}$, providing a reliable and calibrated Q-value estimates of the reference policy, whose values are on the similar scale as the true return value of $\mathcal{D}$ [33]. With $\mathcal{D}$ fixed, we compare the mean estimated return of $\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ sampled from $\mathcal{B}$ against the bootstrapping-error–free ground-truth mean return of $\mathcal{D}$, thereby reducing the accumulated bootstrapping error in the action selection process. Thus, (2) when updating the critic network becomes:
@@
In the bootstrap proposal phase, the future rollouts $s_{t+1}$ are sampled randomly from $\mathcal{B}$. One can select $s^{\prime}_{t+1}$ by finding the states closest to $s_{t+1}$ within $\mathcal{D}$, to enable more precise comparisons between nearby state–action pairs. However, for implementation simplicity, the current $s^{\prime}_{t+1}$ is uniformly sampled from $\mathcal{D}$.
@@ VI-A Bucket-media simulation
TABLE II:
@@ VI-B DRLR implementation
Figure 10: States-actions pairs for training the Ref policy.
@@ VI-C Sim2real results
In the experiments, the inputs $q_{1},q_{2}]$ are measured in radians with Inertial Measurement Units (IMUs) mounted on the boom and bucket. $L_{a}]$ denotes the forwarding distance of the loader, determined using GNSS antennas mounted on the machine. $\hat{\tau}_{e}]$ is computed based on the pressure sensor readings obtained from both sides of the hydraulic pistons in the boom and bucket hydraulic pump. All the sensors operate at an update rate of $10Hz$. The output $q_{d1},q_{d1},\tau_{d}]$ are from the deployed NNs, while the loader’s forwarding motion is manually controlled by an operator at a random speed. The operator halts the forward motion upon noticing the boom’s lift.
Firstly, a two-sided admittance controller with both position and torque reference is tested. However, due to the high compaction nature of wet sand and stone pile, the downward curl of the bucket causes dramatically large normal forces, the admittance controller fails to track $\tau_{d}$, thus leading the boom and bucket to vibrate during penetrating and unstable outputs from NNs. A penalty for causing such unsafe behavior should be considered in the future reward design. For the sake of safety and stable performance, only a one-sided admittance controller is tested in the following experiments with position reference $q_{d1},q_{d2}]$ and a $\tau_{sat}=800$ N.m to prevent the bucket from getting stuck.
@@ VII Conclusion
In future work, one could improve the action selection module by selecting $s^{\prime}_{t+1}$ by finding the states closest to $s_{t+1}$ within $\mathcal{D}$, by employing Euclidean or Mahalanobis distance, thereby facilitating more precise comparisons between neighboring state–action pairs. One could also consider using Deep Ensembles to quantify the uncertainties in the demonstrations and utilize these uncertainties as priors for SAC entropy. Integrating the concepts of Active Learning and Uncertainty-aware RL into the proposed framework could, in theory, further improve the exploration efficiency.

### v2

@@ IV Methods
To reduce the exploration time wasted in correcting unreliable overestimated Q-values, and in turn improves exploration efficiency, it is crucial for the policy to favor distributions whose Q-values are more stable. This motivates selecting batches with reliable Q-value evaluations when updating both the critic and the policy networks. Prior studies have shown that the Q-value estimates of $Q_{\phi^{\prime}}\left(s_{t+1},a(s_{t+1})\right)$ are only reliable when $(s_{t+1},a(s_{t+1}))$ is from the same distributions as the dataset used to train $\hat{Q}(s_{t},{a}_{t})$ [17, 25]. In our critic networks update process, instead of selecting between $Q_{\phi^{\prime}}\left(s_{t+1},\mu_{\psi}(s_{t+1})\right)$ and $Q_{\phi^{\prime}}\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$, where both $\left(s_{t+1},\mu_{\psi}(s_{t+1})\right)$ and $\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ could be OOD state-action pairs. We propose to select between $Q_{\phi^{\prime}}\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ and $Q_{\phi^{\prime}}\left(s^{\prime}_{t+1},\mu_{\psi}(s^{\prime}_{t+1})\right)$, where $s^{\prime}_{t+1}$ are only sampled from $\mathcal{D}$. This modification ensures $(s^{\prime}_{t+1},\mu_{\psi}(s^{\prime}_{t+1}))$ is always from the same distribution as the $\mathcal{D}$, providing a reliable and calibrated Q-value estimates of the reference policy, whose values are on the similar scale as the true return value of $\mathcal{D}$ [25]. With $\mathcal{D}$ fixed, we compare the mean estimated return of $\left(s_{t+1},\pi_{\theta}(s_{t+1})\right)$ sampled from $\mathcal{B}$ against the bootstrapping-error–free ground-truth mean return of $\mathcal{D}$, thereby reducing the accumulated bootstrapping error in the action selection process. Thus, (2) when updating the critic network becomes:
@@
In the bootstrap proposal phase, the future rollouts $s_{t+1}$ are sampled randomly from $\mathcal{B}$. One can select $s^{\prime}_{t+1}$ by finding the states closest to $s_{t+1}$ within $\mathcal{D}$, to enable more precise comparisons between nearby state–action pairs. However, for implementation simplicity, the current $s^{\prime}_{t+1}$ is uniformly sampled from $\mathcal{D}$. By simple random sampling, the expected sample mean Q-value, $\overline{Q}_{\phi^{\prime}}(s^{\prime},\mu_{\psi}(s^{\prime}))$, from each batch converges to the population mean Q-value of the expert buffer [30]. Therefore, even though the comparison is made across different states, it remains valid because we are comparing the mean Q-values of the distributions induced by the IL policy and the RL policy.
@@ VI-A Bucket-media simulation
TABLE II: Domain randomization parameters and their sampling ranges.
@@ VI-B DRLR implementation
Figure 10: States-actions pairs for training the Ref policy. Each curve represents the data recorded in one bucket loading demonstration.
@@ VI-C Sim2real results
In the experiments, the inputs $q_{1},q_{2}]$ are measured in radians with Inertial Measurement Units (IMUs) mounted on the boom and bucket. $L_{a}]$ denotes the forwarding distance of the loader, determined using GNSS antennas mounted on the machine. $\hat{\tau}_{e}]$ is computed based on the pressure sensor readings obtained from both sides of the hydraulic pistons in the boom and bucket hydraulic pump. All the sensors operate at an update rate of $10Hz$. The output $q_{d1},q_{d2},\tau_{d}]$ are from the deployed NNs, while the loader’s forwarding motion is manually controlled by an operator at a random speed. The operator halts the forward motion upon noticing the boom’s lift.
Firstly, a two-sided admittance controller with both position and torque reference is tested. However, due to the high compaction nature of wet sand and stone pile, the downward curl of the bucket causes dramatically large normal forces, the admittance controller fails to track $\tau_{d}$, thus leading the boom and bucket to vibrate during penetrating and unstable outputs from deployed actor network. These unstable NNs outputs could result from a state distribution shift, caused by the large normal forces during interacting with compacted material. In the simulation environment, such compaction effect is not accurately modeled, as the material pile is simulated using discrete particles that lack adhesive or cohesive properties. A penalty for causing such unsafe behavior should be considered in the future reward design. For the sake of safety and stable performance, only a one-sided admittance controller is tested in the following experiments with position reference $q_{d1},q_{d2}]$ and a $\tau_{sat}=800$ N.m to prevent the bucket from getting stuck.
@@ VII Conclusion
In future work, one could improve the action selection module by selecting $s^{\prime}_{t+1}$ by finding the states closest to $s_{t+1}$ within $\mathcal{D}$, by employing Euclidean or Mahalanobis distance, thereby facilitating more precise comparisons between neighboring state–action pairs. To better demonstrate the advantages of DRLR, it is necessary to compare it against established offline-to-online DRL baselines that explicitly addressed bootstrapping errors, such as CAL-QL, RLPD, and WSRL [25, 4, 42].
Moreover, one could also consider using Deep Ensembles to quantify the uncertainties in the demonstrations and utilize these uncertainties as priors for SAC entropy. Integrating the concepts of Active Learning and Uncertainty-aware RL into the proposed framework could further improve the exploration efficiency.